A sports car of mass 950 kg (including the crosses
Answer #1

a) mg-mv^2/r =950*9.8-950*22^2/95 =4470N

b) 72(g-v^2/r) =338.778N

0 = m₂(g - v²/r)
0 = m₂g - m₂v²/r
m₂v²/r = m₂g
v = √gr
= √[(9.8m/s²)(95m)]
= 30.51m/s